A car initially at rest starts moving on straight line with constant acceleration and attains speed og 16m/s in time t s. Average speed of the car in this time interval is

Grade:11

1 Answers

Vikas TU
7 years ago
Re – Posting the answer as asked twice:

From first law of equation, v = u + at 16 = 0 + at a = 16/t m/s^2 Distance covered in t seconds, s = 0 + 0.5*16/t * t^2 => 8t meter. Average speed = tota distance / total time => 8t/t = > 8 m/s.